OpenStudy (anonymous):

Lewis wants to make a mixture of peanuts and candy corn. He wants to have 40% candy corn and 60% peanuts. If he has 12 ounces of peanuts, how many ounces of candy corn should he put in his mix? Show how you arrived at your answer.

OpenStudy (anonymous):

@iGreen, @Luigi0210

OpenStudy (anonymous):

@TeenWolfGirl

OpenStudy (anonymous):

sorry idk

OpenStudy (anonymous):

I think that it is 8 ounces but I dont know how to explain how I got the correct answer

OpenStudy (anonymous):

@iGreen can you help me?

OpenStudy (anonymous):

The answer is 8.

OpenStudy (igreen):

40%/60% = 0.40/0.60 = 2/3 We set up a proportion: \(\dfrac{2}{3} = \dfrac{x}{12}\) Cross multiply: \(3x = 24\) Divide 3 to both sides: \(x = 8\) You are correct! :)

OpenStudy (anonymous):

Okay, but how did you get that answer?

OpenStudy (anonymous):

thank you @iGreen